OpenStudy (anonymous):

Given points A(3, 2) and B(9, 1). Answer the following question concerning using the Pythagorean Theorem to find AB. What is this value of the horizontal leg? 1) 3 2) 1 3) 12 4) 6 ???

OpenStudy (anonymous):

Seems to be a poorly worded question. Is this the question word for word?

OpenStudy (anonymous):

i understand what this is asking

OpenStudy (anonymous):

horizontal leg is 6

OpenStudy (anonymous):

Yeah this is the question word for word .

OpenStudy (anonymous):

The value of the horizontal leg is just the difference between their x values. So x2-x1 would be it. 9-3=6

OpenStudy (anonymous):

from 3 to 9 is 6 to the right

OpenStudy (anonymous):

vertical leg is 1

OpenStudy (anonymous):

and so the hypotenuse, which is the distance between the two points is \[\sqrt{6^2+1^2}=\sqrt{37}\]